# Sneak Path Current Modeling in Memristor Crossbar Arrays for Analog In-Memory Computing

### v1

@@ Abstract
Memristor crossbar arrays have emerged as a key component for next-generation non-volatile memories, artificial neural networks, and analog in-memory computing (IMC) systems. By minimizing data transfer between the processor and memory, they offer substantial energy savings. However, a major design challenge in memristor crossbar arrays is the presence of sneak path currents, which degrade electrical performance, reduce noise margins, and limit reliable operations. This work presents a closed-form analytical framework based on IMEC A14 (1.4 nm) Technology for accurately estimating sneak path currents in memristor crossbar arrays. The proposed model captures the interdependence of key design parameters in memristor crossbar arrays, including array size, ON/OFF ratio of memristors, read voltage, and interconnect conditions, through mathematically derived relationships. It supports various practical configurations, such as different data patterns and connection strategies, enabling rapid and comprehensive sneak path current modeling. The sensitivity analysis includes how design parameters influence sneak path current and noise margin loss, underscoring the trade-offs involved in scaling crossbar arrays. Validation through SPICE simulations shows that the model achieves an error of less than 10.9% while being up to 4784 times faster than full circuit simulations. This analytical framework offers a powerful tool for quantitative assessment and pre-design/real-time optimization of memristor-based analog in-memory computing (IMC) architectures.
@@ I Introduction
MEMRISTOR, widely recognized as the fourth fundamental circuit element, has become central to the development of next-generation non-volatile memory technologies and emerging analog in-memory computing (IMC) systems, addressing the scaling and efficiency challenges of traditional CMOS-based memories [chua2003memristor, pourbakhsh2016sizing, gong2012clock, edstrom2019data, gong2014variation, chen2018viewer, khan2023pawn]. The primary architectural manifestation of memristor-based memory is the high-density passive crossbar array, where each cell is located at the intersection of perpendicular wordlines and bitlines [vourkas2015memristive, shin2011analysis]. This structure provides exceptional storage density since each cell can be scaled to the nanometer regime without requiring a transistor [kim2012functional, jo2009high], thereby offering higher bit-per-area ratios than conventional memories [lastras2018resistive]. Combined with its non-volatility and fast read capability, the memristor crossbar helps bridge the performance gap in modern memory hierarchies [uppaluru2025carbon, 6730965, oli2022stuck, oli2022reliability]. Additionally, memristor plays a crucial role in analog computing by enabling fast and direct hardware implementation of vector matrix multiplication in-memory, such as in a neuromorphic computing system. Memristors can perform computation intrinsically through Ohm’s and Kirchhoff’s laws when placed in crossbar arrays. The output current produced is in the form of weighted summations of input voltage and stored conductance of the memristor. This characteristic of analog IMC results in highly parallel, energy-efficient acceleration for signal processing and Artificial Intelligence (AI) tasks, as it removes the need for frequent data transfer between memory and processors [hu2018memristor].
@@
However, a key challenge in fully realizing the advantages of passive memristor crossbar arrays is the sneak path current [vontobel2009writing]. Sneak paths are unwanted parallel current routes that occur during read operations, diverting current away from the target cell. In unoptimized crossbars, sneak path currents can dominate the total current by exceeding 50%, leading to substantial read errors and increased power consumption [demin2020sneak]. The absence of gating elements and the fully connected structure of the crossbar cause currents to flow preferentially through low-resistance cells [cassuto2013sneak], inevitably creating a parallel resistance network that distorts readouts and causes high-resistance state (HRS) to be misinterpreted as low-resistance state (LRS) [gul2019addressing]. Moreover, the magnitude and distribution of these parasitic currents are highly data-dependent, varying dynamically with the stored information [tang2018comprehensive] or data patterns. Analyzing sneak path currents during the pre-design and ongoing design phases is crucial, as their data-dependent variability poses significant reliability challenges [liang2010size]. Without precise estimation of these parasitic currents, it becomes difficult to ensure stable performance in terms of noise margin, functional yield, and maximum achievable array size [chen2013comprehensive].
@@
Some research has focused on mitigating this issue through various strategies. Architectural approaches often incorporate gating elements, such as diodes, transistors, or additional memristors, to suppress unwanted current flow, although at the cost of reduced array density. Alternatively, gateless techniques leverage multi-point read schemes, coding strategies, or logical operations to achieve sneak path free reading [zidan2013memristor, joshi2021sneak, manem2010design, linn2010complementary, shi2020research, kannan2013sneak]. However, the above studies only limit to the individual aspects such as data patterns, connection configurations, read voltages, or array sizes, rather than capture the combined influence of all key parameters. Consequently, a unified analytical framework that comprehensively models sneak path behavior across multiple interacting factors remains largely unexplored. In addition, recent Computer-Aided Design (CAD) frameworks such as CiMLoop [andrulis2024cimloop], DNN+NeuroSim [luo2019mlp+], MNSIM [zhu2023mnsim], and CrossSim [maram2023deep] have advanced the modeling and co-design of IMC architectures. For example, CiMLoop enables full-stack exploration from device to architecture; DNN+NeuroSim provides circuit to algorithm benchmarking for SRAM and emerging non-volatile memories; MNSIM supports hierarchical modeling of digital and analog IMC architectures; and CrossSim simulates ReRAM-based neural training under realistic device nonidealities. However, these tools neither consider sneak path effects nor employ purely passive memristor arrays; instead, they depend on selector or switching devices to mitigate interference, which substantially limits memory density. This highlights the need for a comprehensive analytical framework dedicated to modeling sneak path phenomena in high-density passive crossbar arrays.
@@ II-A Design Considerations
The foundation of this analysis is the nanoscale crossbar memristor model presented in [robinett2007demultiplexers] based on HP Technology, where the device current is defined in the following equation:
@@
In this configuration, a TiO2 layer for a total thickness of the device, ${h}\leq 10\,\text{nm}$, is sandwiched between two platinum electrodes. The oxide layer has two constituent layers: a doped region containing oxygen vacancies and an undoped region composed of pure TiO2, as shown in Fig. 1 (a). This work is conducted using a parameter value of $\alpha$ = 3 [strukov2008missing] and current versus voltage characteristic for the memristor is shown in Fig. 1 (d) [robinett2007demultiplexers].
@@
Figure 1: (a) Structure of the linear dopant drift memristor model. h represents the total thickness of the device. w represents the thickness of the doped region; (d) Current versus voltage characteristic for the memristor [robinett2007demultiplexers]. A $3\times 3$ memristive crossbar, where Vdd = V2 is the voltage applied in row 2, and current sensed from column 2: (b) the crossbar without sneak paths where the desired current path is marked in green line, (c) presence of a sneak path which is marked in red line, (e) FRC connection strategy with ‘All Ones’ (LRS) data pattern, (f) GRC connection strategy with ‘All Zeros’ (HRS) data pattern. Here, in GRC connection strategy, shorted row nodes are indicated as "nrs" and shorted column nodes are indicated as "ncs".
@@
The current method presented in this paper employs the constant Kon and Koff as the ON and OFF states, respectively. To comprehensively assess the influence of sneak path currents, two representative data patterns are simulated: the ‘All Zeros’ pattern (best case scenario, minimum sneak path current effect, all the memristors are in the OFF state and the sneak paths are made of OFF resistances in series); and the ‘All Ones’ pattern (worst case scenario, maximum sneak path current effect). Grounded unselected rows and columns has been explored as a method to mitigate sneak path currents. Grounded rows and columns of unselected cells in an array provides alternating paths for leakage currents to flow instead of the sensing circuit. However, part of the sneak path current still flows through the sense circuit. Four terminal biasing configurations are analyzed: Floating Rows and Columns (FRC), Grounded Rows with Floating Columns (GRFC), Floating Rows with Grounded Columns (FRGC), and Grounded Rows and Columns (GRC), to examine the impact of different connection strategies on the overall array behavior. The grounded rows and columns are implemented using $0.001\,\Omega$ resistors (near zero), with the corresponding nodes labeled as node row shorted and node column shorted, respectively.
@@
Fig. 1 presents an abstract depiction of the sneak path phenomenon in the crossbar array, along with schematic representations to illustrate the different combinations of data patterns and connection strategies. The device at the center of the array is the “target device," and the devices sharing a line with the target are called “half-selected devices". To read the “target device", the selected row is pulled up to Vdd and the selected column is grounded through a resistor. Sneak path occurs when current flows from Vdd to ground through unselected devices. Therefore, it affects the accuracy of the read operation by degrading the voltage and current delivered to the device. The crossbar interconnect network is modeled at the unit-cell level to accurately capture parasitic effects, incorporating line resistance, stray capacitance to ground, and fringing capacitance between adjacent lines. In this work, the interconnect design dimensions follow the IMEC A14 (1.4 nm) technology node, offering concrete guidance for both academic researchers and industry engineers [pei2025interconnect, pei2024ultra, liu2023cfet, pei2023technology, liu2024future, liu2023cfetted]. For the IMEC A14 interconnect, metal layers M1 and M2 are primarily employed for internal interconnect within the cell. Layers M1 and M3 exhibit identical geometries, as do M4 and M5. However, M3 can be used for local interconnect within subarrays, and M4 and M5 are for intermediate-level interconnect. Metal layers M6 and above maintain consistent geometries for intermediate- and global-level interconnects. Consequently, layers M3, M5, and M6 are selected to represent interconnects spanning from local to global scales and are investigated. Fig. 2 (a) presents the utility of metal layers (M1~M12) for different interconnects. The geometry of the Cu interconnect and the pitch (P) is shown in Fig. 2 (b). Here, W, T represent the width and thickness of Cu. The thickness is represented as the product of the width (W) and aspect ratio (ar), and the pitch (P) represents the distance between the centers of two adjacent interconnects. The unit-cell memristor array is shown in Fig. 2 (c) with the corresponding values of W, T, and P. The selected interconnect metal layers provide different parasitic resistance Rline and capacitance Cline. The corresponding ranges of interconnect parameters are listed in Table I.
@@
Metal W(nm) P(nm) T(nm) $R_{\text{line}}$($\Omega$) $C_{\text{line}}$(fF) M3 16 28 49 3.122 3.871e-3 M5 16 28 28 5.869 2.871e-3 M6 40 80 80 0.7396 1.020e-2
@@
Figure 2: (a) Metal layers M1~M12 are used for interconnect within the cell; (b) Sketch of Cu for interconnect material; (c) Top view of a unit-cell memristor crossbar array for M5/M4 with IMEC A14.
@@ III-B Noise Margin Analysis
For the analysis of the sneak path effect on the noise margin of the memristor array, a normalized parameter is defined [zidan2013memristor]:
@@ III-C Interconnect Resistance Effect on Sneak Path Current
Figure 5: (a) Sneak path current and (b) noise margin of the target cell versus array size for the interconnect layers using the IMEC A14 technology node. The array adopts ‘All Ones’ data pattern with FRC connection strategy. Kon is 1e-7. Vdd is 3 V.
Figure 6: Relative percentage change in sneak path current and noise margin as Kon and Vdd are varied. The array adopts ‘All Ones’ data pattern with FRC connection strategy using the IMEC A14 technology node with Cu M3 interconnect.
@@ IV Comparison with State-of-The-Art
Device Type [zidan2013memristor] [manem2010design] [joshi2021sneak] [datta2022unlocking] Current
@@
Read voltage consideration $\surd$ × x $\surd$ $\surd$
@@ Acknowledgment
[Uncaptioned image] Shah Zayed Riam received the B.Sc. and M.S. degrees in electrical engineering from Rajshahi University of Engineering and Technology (RUET), Bangladesh, and the University of Texas at Tyler, USA, respectively, in 2022 and 2024. He is currently pursuing a Ph.D. degree at the University of Alabama, Tuscaloosa, AL, USA. His research interests include memory design, flexible electronics, and electrochemical biosensors. His previous works received a runner-up award at WiSe/YP BIP 2023 and a nomination for the Best Paper Award at IEEE FLEPS 2023, Boston, MA, USA.
[Uncaptioned image] Zhenlin Pei (Graduate Student Member, IEEE) holds the M.S. and Ph.D. degrees in electrical engineering from Columbia University and the University of Texas at Arlington. During his Ph.D., he collaborated with the Interuniversity Microelectronics Centre to develop the Cacti++ framework, which bridges gaps in rapid EDA-based co-design and co-optimization from the transistor to the system level. He is a postdoctoral fellow in the department of electrical & computer engineering at the University of Alabama. He was a senior design engineer in the IP group for tapeout at Cadence for four years. Research interests include energy-efficient computing from hardware to software, with applications in AI through CAD/EDA & DTCO/STCO, emerging interconnects/beyond-CMOS/memory technologies, AI hardware with privacy, and neuromorphic systems.
[Uncaptioned image] Kyle Mooney (Graduate Student Member, IEEE) received his B.S. degree in computer engineering from the University of South Alabama in 2023 and is currently pursuing a Ph.D. in Electrical Engineering from the University of Alabama. His research interests include memory design, edge computing, and artificial intelligence.
[Uncaptioned image] Chenyun Pan (Senior Member, IEEE) received a B.S. in Microelectronics from Shanghai Jiao Tong University and a Ph.D. in ECE from Georgia Institute of Technology. He is an Associate Professor at the University of Texas at Arlington, and his research focuses on energy-efficient Boolean and non-Boolean computing systems using emerging technologies. He has published over 80 IEEE papers and received several awards, including the Research Spotlight Award from Georgia Institute of Technology, and an early career research award from the U.S. Department of Energy.
[Uncaptioned image] Na Gong (Senior Member, IEEE) received the Ph.D. degree in computer science and engineering from the State University of New York, Buffalo, in 2013. Currently, Dr. Gong is a professor in the Department of Electrical and Computer Engineering at the University of Alabama. Her research interests include power-efficient computing circuits and systems, memory optimization, AI hardware, and hardware privacy. She is the recipient of the best paper nomination from ISVLSI’19, best paper award from EIT’16, best paper nominations from ISQED’16 and ISLPED’16.
[Uncaptioned image] Jinhui Wang (Senior Member, IEEE) is currently a Full Professor and Larry Drummond Endowed Chair with the Department of Electrical and Computer Engineering at the University of Alabama, Tuscaloosa, AL, USA. His research interests include: (1) VLSI System, Digital and Mixed-Signal Integrated Circuit (IC) Design, 3D and 2.5D IC Design, and Emerging Memory; (2) AI Hardware Design, Post/Beyond CMOS Device, such as Memristors, Based Neuromorphic Computing System; and (3) Post/Beyond CMOS Devices Enabled Cybersecurity and Internet of Things (IoT) Systems. He has published over 200 refereed journal/conference papers and book chapters as well as 31 patents in the area of emerging semiconductor technologies. His previous work has received the Best Paper Award/Nomination at DATE 2021, ISVLSI 2019, ISLPED 2016, ISQED 2016, and EIT 2016.

### v3

@@ Abstract
Memristor crossbar arrays have emerged as a key component for next-generation non-volatile memories, artificial neural networks, and analog in-memory computing (IMC) systems. By minimizing data transfer between the processor and memory, they offer substantial energy savings. However, a major design challenge in memristor crossbar arrays is the presence of sneak path currents, which degrade electrical performance, reduce noise margins, and limit reliable operations. This work presents a closed-form analytical framework based on 1.4nm technology for accurately estimating sneak path currents in memristor crossbar arrays. The proposed model captures the interdependence of key design parameters in memristor crossbar arrays, including array size, ON/OFF ratio of memristors, read voltage, and interconnect conditions, through mathematically derived relationships. It supports various practical configurations, such as different data patterns and connection strategies, enabling rapid and comprehensive sneak path current modeling. The sensitivity analysis includes how design parameters influence sneak path current and noise margin loss, underscoring the trade-offs involved in scaling crossbar arrays. Validation through SPICE simulations shows that the model achieves an error of less than 10.9% while being up to 4784 times faster than full circuit simulations. This analytical framework offers a powerful tool for quantitative assessment and pre-design/real-time optimization of memristor-based analog in-memory computing (IMC) architectures.
@@ I Introduction
MEMRISTOR, widely recognized as the fourth fundamental circuit element, has become central to the development of next-generation non-volatile memory technologies and emerging analog in-memory computing (IMC) systems, addressing the scaling and efficiency challenges of traditional CMOS-based memories [5, 33, 9, 8, 11, 4, 17]. The primary architectural manifestation of memristor-based memory is the high-density passive crossbar array, where each cell is located at the intersection of perpendicular wordlines and bitlines [41, 36]. This structure provides exceptional storage density since each cell can be scaled to the nanometer regime without requiring a transistor [18, 14], thereby offering higher bit-per-area ratios than conventional memories [19]. Combined with its non-volatility and fast read capability, the memristor crossbar helps bridge the performance gap in modern memory hierarchies [39, 10, 28, 29]. Additionally, memristor plays a crucial role in analog computing by enabling fast and direct hardware implementation of vector matrix multiplication in-memory, such as in a neuromorphic computing system. Memristors can perform computation intrinsically through Ohm’s and Kirchhoff’s laws when placed in crossbar arrays. The output current produced is in the form of weighted summations of input voltage and stored conductance of the memristor. This characteristic of analog IMC results in highly parallel, energy-efficient acceleration for signal processing and Artificial Intelligence (AI) tasks, as it removes the need for frequent data transfer between memory and processors [13].
@@
However, a key challenge in fully realizing the advantages of passive memristor crossbar arrays is the sneak path current [40]. Sneak paths are unwanted parallel current routes that occur during read operations, diverting current away from the target cell. In unoptimized crossbars, sneak path currents can dominate the total current by exceeding 50%, leading to substantial read errors and increased power consumption [7]. The absence of gating elements and the fully connected structure of the crossbar cause currents to flow preferentially through low-resistance cells [2], inevitably creating a parallel resistance network that distorts readouts and causes high-resistance state (HRS) to be misinterpreted as low-resistance state (LRS) [12]. Moreover, the magnitude and distribution of these parasitic currents are highly data-dependent, varying dynamically with the stored information [38] or data patterns. Analyzing sneak path currents during the pre-design and ongoing design phases is crucial, as their data-dependent variability poses significant reliability challenges [20]. Without precise estimation of these parasitic currents, it becomes difficult to ensure stable performance in terms of noise margin, functional yield, and maximum achievable array size [3].
@@
Some research has focused on mitigating this issue through various strategies. Architectural approaches often incorporate gating elements, such as diodes, transistors, or additional memristors, to suppress unwanted current flow, although at the cost of reduced array density. Alternatively, gateless techniques leverage multi-point read schemes, coding strategies, or logical operations to achieve sneak path free reading [43, 15, 26, 21, 35, 16]. However, the above studies only limit to the individual aspects such as data patterns, connection configurations, read voltages, or array sizes, rather than capture the combined influence of all key parameters. Consequently, a unified analytical framework that comprehensively models sneak path behavior across multiple interacting factors remains largely unexplored. In addition, recent Computer-Aided Design (CAD) frameworks such as CiMLoop [1], DNN+NeuroSim [25], MNSIM [42], and CrossSim [27] have advanced the modeling and co-design of IMC architectures. For example, CiMLoop enables full-stack exploration from device to architecture; DNN+NeuroSim provides circuit to algorithm benchmarking for SRAM and emerging non-volatile memories; MNSIM supports hierarchical modeling of digital and analog IMC architectures; and CrossSim simulates ReRAM-based neural training under realistic device nonidealities. However, these tools neither consider sneak path effects nor employ purely passive memristor arrays; instead, they depend on selector or switching devices to mitigate interference, which substantially limits memory density. This highlights the need for a comprehensive analytical framework dedicated to modeling sneak path phenomena in high-density passive crossbar arrays.
@@ II-A Design Considerations
The foundation of this analysis is the nanoscale crossbar memristor model presented in [34] based on HP Technology, where the device current is defined in the following equation:
@@
In this configuration, a TiO2 layer for a total thickness of the device, ${h}\leq 10\,\text{nm}$, is sandwiched between two platinum electrodes. The oxide layer has two constituent layers: a doped region containing oxygen vacancies and an undoped region composed of pure TiO2, as shown in Fig. 1 (a). This work is conducted using a parameter value of $\alpha$ = 3 [37] and current versus voltage characteristic for the memristor is shown in Fig. 1 (d) [34].
@@
Figure 1: (a) Structure of the linear dopant drift memristor model. h represents the total thickness of the device. w represents the thickness of the doped region; (d) Current versus voltage characteristic for the memristor [34]. A $3\times 3$ memristive crossbar, where Vdd = V2 is the voltage applied in row 2, and current sensed from column 2: (b) the crossbar without sneak paths where the desired current path is marked in green line, (c) presence of a sneak path which is marked in red line, (e) FRC connection strategy with ‘All Ones’ (LRS) data pattern, (f) GRC connection strategy with ‘All Zeros’ (HRS) data pattern. Here, in GRC connection strategy, shorted row nodes are indicated as "nrs" and shorted column nodes are indicated as "ncs".
@@
The current method presented in this paper employs the constant Kon and Koff as the ON and OFF states, respectively. To comprehensively assess the influence of sneak path currents, two representative data patterns are simulated: the ‘All Zeros’ pattern (best case scenario, minimum sneak path current effect, all the memristors are in the OFF state and the sneak paths are made of OFF resistances in series); and the ‘All Ones’ pattern (worst case scenario, maximum sneak path current effect). Grounded unselected rows and columns has been explored as a method to mitigate sneak path currents. Grounded rows and columns of unselected cells in an array provides alternating paths for leakage currents to flow instead of the sensing circuit. However, part of the sneak path current still flows through the sense circuit. Four terminal biasing configurations are analyzed: Floating Rows and Columns (FRC), Grounded Rows with Floating Columns (GRFC), Floating Rows with Grounded Columns (FRGC), and Grounded Rows and Columns (GRC), to examine the impact of different connection strategies on the overall array behavior. The grounded rows and columns are implemented using $0.01\,\Omega$ resistors (near zero), with the corresponding nodes labeled as node row shorted and node column shorted, respectively.
@@
Fig. 1 presents an abstract depiction of the sneak path phenomenon in the crossbar array, along with schematic representations to illustrate the different combinations of data patterns and connection strategies. The device at the center of the array is the “target device," and the devices sharing a line with the target are called “half-selected devices". To read the “target device", the selected row is pulled up to Vdd and the selected column is grounded through a resistor. Sneak path occurs when current flows from Vdd to ground through unselected devices. Therefore, it affects the accuracy of the read operation by degrading the voltage and current delivered to the device. The crossbar interconnect network is modeled at the unit-cell level to accurately capture parasitic effects, incorporating line resistance, stray capacitance to ground, and fringing capacitance between adjacent lines. In this work, the interconnect design dimensions follow the 1.4nm technology node, offering concrete guidance for both academic researchers and industry engineers [31, 30, 23, 32, 22, 24]. Generally, three interconnect layers M3, M5, and M6 with different geometries are chosen to represent interconnects at local, intermediate, and global levels, respectively, as illustrated in Fig. 2. M3 can be used for local interconnects. M5 is for the intermediate-level interconnects. M6 and above are for intermediate- and global-level interconnects. Here, the interconnect geometry is characterized by the width (W), thickness (T), and pitch (P), as illustrated in Fig. 2 (a). The aspect ratio (AR) is defined as the ratio of T to W, while P denotes the center-to-center spacing between adjacent interconnects. The unit-cell memristor array, annotated with the corresponding values of W and P, is depicted in Fig. 2 (b). Fig. 2 (c) presents the utility of metal layers for different interconnects. The selected interconnect metal layers provide different parasitic resistance Rline and capacitance Cline. The corresponding values of interconnect parameters are listed in Table I.
@@
Metal W(nm) P(nm) T(nm) Rline($\Omega$) Cline(fF) M3 16 28 49 3.122 3.871e-3 M5 16 28 28 5.869 2.871e-3 M6 40 80 80 0.7396 1.020e-2
@@
Figure 2: (a) Cu interconnect geometry, with barrier thickness taken into account; (b) Top view of the unit-cell memristor crossbar array for M5 with 1.4nm technology; (c) Metal layers are used for interconnects within cells.
@@ III-B Noise Margin Analysis
For the analysis of the sneak path effect on the noise margin of the memristor array, a normalized parameter is defined [43]:
@@ III-C Interconnect Resistance Effect on Sneak Path Current
Figure 5: (a) Sneak path current and (b) noise margin of the target cell versus array size for the interconnect layers using the 1.4nm technology. The array adopts ‘All Ones’ data pattern with FRC connection strategy. Kon is 1e-7. Vdd is 3 V.
Figure 6: Relative percentage change in sneak path current and noise margin as Kon and Vdd are varied. The array adopts ‘All Ones’ data pattern with FRC connection strategy using the 1.4nm technology node with Cu M3 interconnect.
@@ IV Comparison with State-of-The-Art
Device Type [43] [26] [15] [6] Current
@@
Read voltage consideration $\surd$ × × $\surd$ $\surd$